I have a lot 100 feet by 80 feet. A fence is going around the lot. How many feet of fencing will I need?
Answer by TimothyLamb(4379)

2L + 2w = p
2*100 + 2*80 = p
200 + 160 = p
---
p = 360 ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php